# 0-supple.tex

### v1

@@ 3.1 Distributed Tensor PCA for Heterogeneous Tensors
Then we send $\widehat{U}_{j,\ell}$ to the central machine and aggregate the projection matrices $\widehat{U}_{j,\ell}\widehat{U}_{j,\ell}^{\top}$ to compute a global estimator $\widehat{U}_{j}$. To estimate $\mathrm{Col}(V_{j,\ell})$, we further send $\widehat{U}_{j}$ back to each machine and compute the top singular vectors of $\big{(}I_{p_{j}}-\widehat{U}_{j}\widehat{U}_{j}^{\top}\big{)}\widetilde{M}_{j,\ell}$, the projection of $\widetilde{M}_{j,\ell}$ onto the orthogonal space of $\mathrm{Col}(\widehat{U}_{j})$, whose top singular subspace provides a local estimator for $\mathrm{Col}(V_{j,\ell})$.
@@ 3.2 Theoretical Guarantee
Assume that there exist constants $C_{1},c_{1},C_{2}$ such that, with probability at least $1-C_{1}e^{-c_{1}p}$, $\big{\|}\widehat{U}_{j,\ell}^{(0)}\widehat{U}_{j,\ell}^{(0)\top}+\widehat{V}_{j,\ell}^{(0)}\widehat{V}_{j,\ell}^{(0)\top}-U_{j}U_{j}^{\top}-V_{j,\ell}V_{j,\ell}^{\top}\big{\|}_{2}\leq C_{2}\sqrt{p}\sigma\lambda_{\min}^{-1}$ for all $j,\ell$. If $p_{j}\asymp p$ for all $j$, $L\lesssim p^{c_{3}}$ for some $c_{3}>0$, $p\gtrsim r^{J-1}$, $\kappa_{0}=O(1)$, $\kappa=O(1)$, and $\min(\Delta,\lambda_{\min})/\sigma\gtrsim\sqrt{pr}$, then we have
$\sup_{j}\rho\big{(}\widehat{U}_{j},U_{j}\big{)}=O_{\mathbb{P}}\left(\sqrt{\frac{pr}{L}}\frac{\sigma}{\Delta}+\frac{pr\sigma^{2}}{\Delta^{2}}\right),$ (10)
@@
$\sup_{j,\ell}\rho\big{(}\widehat{V}_{j,\ell},V_{j,\ell}\big{)}=O_{\mathbb{P}}\left(\frac{\sqrt{pr_{V}}\sigma(1+\sqrt{r/L})}{\lambda_{\min}}+\frac{\sqrt{r_{V}}pr\sigma^{2}}{\lambda_{\min}^{2}}\right),$ (11)
where $\widehat{U}_{j}$ and $\widehat{V}_{j,\ell}$’s are the output of Algorithm 2.
@@ 4 Knowledge Transfer in Distributed Tensor PCA
In this section, we explore the task of transferring knowledge from source locations to a target location within a heterogeneous setting. Knowledge Transfer seeks to enhance learning performance at a target site by leveraging insights from related source tasks. To illustrate our approach clearly, we concentrate on transferring knowledge between a single source dataset and a target dataset. However, the transfer of knowledge across multiple tasks can be managed by integrating Algorithm 2 with Algorithm 3.
@@ 4.1 Transferred Tensor PCA
10:     Compute $\widehat{V}_{j,t}=\mathrm{svd}_{r_{j,V,\ell}}\Big{[}\big{(}I_{p_{j}}-\widehat{U}_{j}\widehat{U}_{j}^{\top}\big{)}\widetilde{M}_{j,t}\Big{]}$;
@@
$\mathcal{T}_{s}^{*}$ $=\mathcal{G}_{s}\times_{1}[U_{1}\;V_{1,s}]\times_{2}[U_{2}\;V_{2,s}]\cdots\times_{J}[U_{J}\;V_{J,s}],\quad U_{j}\in\mathbb{O}^{p_{j}\times r_{j,U}},V_{j,s}\in\mathbb{O}^{p_{j}\times r_{j,V,s}},\mathcal{G}_{s}\in\mathbb{R}^{r_{1,s}\times r_{2,s}\cdots\times r_{J,s}},$
$\mathcal{T}_{t}^{*}$ $=\mathcal{G}_{t}\times_{1}[U_{1}\;V_{1,t}]\times_{2}[U_{2}\;V_{2,t}]\cdots\times_{J}[U_{J}\;V_{J,t}],\quad U_{j}\in\mathbb{O}^{p_{j}\times r_{j,U}},V_{j,t}\in\mathbb{O}^{p_{j}\times r_{j,V,t}},\mathcal{G}_{t}\in\mathbb{R}^{r_{1,t}\times r_{2,t}\cdots\times r_{J,t}},$
with $U_{j}^{\top}V_{j,\ell}=0$ and $r_{j,\ell}=r_{j,U}+r_{j,V,\ell}$ for $j\in[J]$ and $\ell=s,t$. In other words, we assume the source and target tensors share a common top-$r_{j,U}$ singular subspace spanned by $U_{j}$, but either task can have different individual components $V_{j,\ell}$. The goal is to estimate the singular subspaces spanned by $U_{j}$ and $V_{j,t}$ of the target tensor $\mathcal{T}_{t}$. Meanwhile, we assume the noise $\mathcal{Z}_{\ell}$ has i.i.d. $\mathcal{N}(0,\sigma_{\ell}^{2})$ entries for $\ell=s,t$, where the noise levels $\sigma_{s}$ and $\sigma_{t}$ are allowed to be different.
To achieve knowledge transfer between $\mathcal{T}_{s}$ and $\mathcal{T}_{t}$, we propose Algorithm 3, which is carefully designed for dealing with the heterogeneity in the transfer setting. Different from Algorithm 2 who treats all tensors equally, Algorithm 3 aggregates the local estimators $\widehat{U}_{j,s}$ and $\widehat{U}_{j,t}$ through a weighted average. The weights $w_{s}$ and $w_{t}$ are designed to optimally balance the contributions from the source and target tensors, respectively, accounting for the potential heterogeneity in their noise levels $\sigma_{s}^{2}$ and $\sigma_{t}^{2}$. The choice for the weights will be specified in the next section.
@@ 4.2 Theoretical Guarantee
Assume that there exist constants $C_{1},c_{1},C_{2}$ such that, with probability at least $1-C_{1}e^{-c_{1}p}$, $\big{\|}\widehat{U}_{j,\ell}^{(0)}\widehat{U}_{j,\ell}^{(0)\top}+\widehat{V}_{j,\ell}^{(0)}\widehat{V}_{j,\ell}^{(0)\top}-U_{j}U_{j}^{\top}-V_{j,\ell}V_{j,\ell}^{\top}\big{\|}_{2}\leq C_{2}\sqrt{p}\sigma_{\ell}\lambda_{\min}^{-1}$ for all $j\in[J],\ell\in\{s,t\}$. Assume $p_{j}\asymp p$ for all $j$, $p\gtrsim r^{J-1}$, $\kappa_{0}=O(1)$, $\kappa=O(1)$, and $\min(\Delta,\lambda_{\min})/\max(\sigma_{t},\sigma_{s})\gtrsim\sqrt{pr}$. We have
$\sup_{j}\left\lVert\widehat{U}_{j}\widehat{U}_{j}^{\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}=$ $~{}O_{\mathbb{P}}\left(\frac{\sqrt{pr}\sqrt{w_{s}^{2}\sigma_{s}^{2}+w_{t}^{2}\sigma_{t}^{2}}}{\Delta}\right),$ (12)
$\sup_{j}\left\lVert\widehat{V}_{j,t}\widehat{V}_{j,t}^{\top}-V_{j,t}V_{j,t}^{\top}\right\rVert_{\rm F}=$ $~{}O_{\mathbb{P}}\left(\frac{\sqrt{pr_{V}}}{\lambda_{\min}}\Big{(}\sigma_{t}+\sqrt{r}\sqrt{w_{s}^{2}\sigma^{2}_{s}+w_{t}^{2}\sigma_{t}^{2}}\Big{)}\right),$ (13)
where $\widehat{U}_{j},\widehat{V}_{j,t}$’s are the outputs of Algorithm 3.
@@
where $\overline{\sigma}^{2}=1/(\sigma_{s}^{-2}+\sigma_{t}^{-2})$. In practice, we can estimate $\sigma_{s}$ and $\sigma_{t}$ by
$\widehat{\sigma}_{\ell}=\Big{\|}\mathcal{T}_{\ell}-\mathcal{T}_{\ell}\times_{1}\big{[}\widehat{U}_{1,\ell}\;\widehat{V}_{1,\ell}\big{]}\times_{2}\big{[}\widehat{U}_{2,\ell}\;\widehat{V}_{2,\ell}\big{]}\cdots\times_{J}\big{[}\widehat{U}_{J,\ell}\;\widehat{V}_{J,\ell}\big{]}\Big{\|}_{\rm F}/\sqrt{p_{1}p_{2}\cdots p_{J}},\quad\ell=s,t.$ (14)
@@ 5 Numerical Study
In this section, we conduct numerical studies to verify the theoretical properties and evaluate the empirical performance of our proposed distributed tensor PCA algorithms. Section 5.1 presents simulation studies for both the homogeneous and heterogeneous settings. Section 5.2 then illustrates the usefulness of our algorithms on a real dataset of protein structure.

### v3

@@ 3.1 Distributed Tensor PCA for Heterogeneous Tensors
We then send $\widehat{U}_{j,\ell}$ to the central machine and aggregate the projection matrices $\widehat{U}_{j,\ell}\widehat{U}_{j,\ell}^{\top}$ to compute a global estimator $\widehat{U}_{j}$. To estimate $\mathrm{Col}(V_{j,\ell})$, we send $\widehat{U}_{j}$ back to each machine and
compute the top singular vectors of $\big{(}I_{p_{j}}-\widehat{U}_{j}\widehat{U}_{j}^{\top}\big{)}\widetilde{M}_{j,\ell}$, the projection of $\widetilde{M}_{j,\ell}$ onto the orthogonal space of $\mathrm{Col}(\widehat{U}_{j})$, whose top singular subspace provides a local estimator for $\mathrm{Col}(V_{j,\ell})$.
@@ 3.2 Theoretical Guarantee
Assume that there exist constants $C_{1},c_{1},C_{2}$ such that, with probability at least $1-C_{1}e^{-c_{1}p}$, $\big{\|}\widehat{U}_{j,\ell}^{(0)}\widehat{U}_{j,\ell}^{(0)\top}+\widehat{V}_{j,\ell}^{(0)}\widehat{V}_{j,\ell}^{(0)\top}-U_{j}U_{j}^{\top}-V_{j,\ell}V_{j,\ell}^{\top}\big{\|}_{2}\leq C_{2}\sqrt{p}\sigma\lambda_{\min}^{-1}$ for all $j,\ell$. If $p_{j}\asymp p$ for all $j$, $L=O(p^{c_{3}})$ for some $c_{3}>0$, $r^{J-1}=O(p)$, $\kappa_{0}=O(1)$, $\kappa=O(1)$, and $\sqrt{pr}=o(\min(\Delta,\lambda_{\min})/\sigma)$, then we have
$\sup_{j}\rho\big{(}\widehat{U}_{j},U_{j}\big{)}\leq\widetilde{C}_{2}\left(\sqrt{\frac{pr}{L}}\frac{\sigma}{\Delta}+\frac{pr\sigma^{2}}{\Delta^{2}}\right),$ (10)
@@
$\sup_{j,\ell}\rho\big{(}\widehat{V}_{j,\ell},V_{j,\ell}\big{)}\leq\widetilde{C}_{2}\left(\frac{\sqrt{pr_{V}}\sigma(1+\sqrt{r/L})}{\lambda_{\min}}+\frac{\sqrt{r_{V}}pr\sigma^{2}}{\lambda_{\min}^{2}}\right),$ (11)
with probability at least $1-\widetilde{C}_{1}e^{-\widetilde{c}_{1}p}$ for some constants $\widetilde{C}_{1}$ , $\widetilde{c}_{1}$, and $\widetilde{C}_{2}$, where $\widehat{U}_{j}$ and $\widehat{V}_{j,\ell}$’s are the output of Algorithm 2.
@@
It is worth noting that, the definition of $\Delta$ is related to our theoretical framework that assumes the common components $U_{j}$ correspond to the large singular values for clear presentation. While extension to settings where $U_{j}$’s are not necessarily the top singular vectors is technically feasible, we leave it to future work as the analysis would involve introducing numerous additional notations and case-specific definitions without providing significant new insights. We anticipate that the error rate will remain the same as Theorem 3.1, requiring only a redefinition of the signal strength $\Delta$. This redefined $\Delta$ still represents the minimum eigenvalue gap between singular values of common and individual components, but writing out its explicit form would be notationally cumbersome.
3.3 Rank Determination
In this section, we provide an estimation method for determining the ranks $r_{j,U}$ and $r_{j,V,\ell}$ in model (7) if they are unknown in practice. Note that in Remark 2, we discuss how to use the existing rank determination methods to obtain an accurate estimate for $r_{j}$ under the homogeneous setting. Similarly, these methods can be applied to estimate the rank of $\widetilde{M}_{j,\ell}$ defined in (9), i.e., $r_{j,U}+r_{j,V,\ell}$. Therefore, we focus on how to obtain a consistent estimate for $r_{j,U}$, assuming that we have already correctly estimated $\widehat{r}_{j,\ell}=r_{j,U}+r_{j,V,\ell}$ for all $\ell$. For each tensor $\mathcal{T}_{\ell}$, compute the left $\widehat{r}_{j,\ell}$ singular vectors of $\widetilde{M}_{j,\ell}$ and denote it by $\widehat{W}_{j,\ell}$. We then compute $\overline{W}_{j}=\frac{1}{L}\sum_{\ell=1}^{L}\widehat{W}_{j,\ell}\widehat{W}_{j,\ell}^{\top}$ and its singular values $\sigma_{1}\big{(}\overline{W}_{j}\big{)}\geq\sigma_{2}\big{(}\overline{W}_{j}\big{)}\geq\dots\geq\sigma_{p_{j}}\big{(}\overline{W}_{j}\big{)}$. Finally, we estimate $r_{j,U}$ as
$\widehat{r}_{j,U}=\max\left\{k\mid\sigma_{k}\big{(}\overline{W}_{j}\big{)}\geq 1-\delta_{0}\right\},$
for some constant $\delta_{0}>0$. The motivation behind $\widehat{r}_{j,U}$ stems from the fact that $\overline{W}_{j}$ is a consistent estimator for the matrix $U_{j}U_{j}^{\top}+\frac{1}{L}\sum_{\ell=1}^{L}V_{j,\ell}V_{j,\ell}^{\top}$. Its eigenvalues exhibit a distinctive pattern: those corresponding to $U_{j}U_{j}^{\top}$ are 1, whereas those corresponding to $\frac{1}{L}\sum_{\ell=1}^{L}V_{j,\ell}V_{j,\ell}^{\top}$ should be less than 1 due to the heterogeneity among $\mathrm{Col}(V_{j,\ell})$ subspaces. To formalize this intuition, we introduce a heterogeneity measure $\eta_{V}:=\min_{j,\ell}\frac{1}{L-1}\sum_{\ell^{\prime}\neq\ell}(1-\left\lVert V_{j,\ell}^{\top}V_{j,\ell^{\prime}}\right\rVert_{2})$, which quantifies the dissimilarity among the individual components $V_{j,\ell}$ across different tensors. The largest singular value of $V_{j,\ell}^{\top}V_{j,\ell^{\prime}}$, i.e., $\left\lVert V_{j,\ell}^{\top}V_{j,\ell^{\prime}}\right\rVert_{2}$, represents the cosine of the smallest principal angle between the subspaces $\mathrm{Col}(V_{j,\ell})$ and $\mathrm{Col}(V_{j,\ell^{\prime}})$. A larger value of $\left\lVert V_{j,\ell}^{\top}V_{j,\ell^{\prime}}\right\rVert_{2}$ indicates a smaller distance between the corresponding principal vectors of these subspaces. Consequently, $\eta_{V}$ can be interpreted as the minimum (over $\ell$) of the average distance between the principal vectors of $\mathrm{Col}(V_{j,\ell})$ and $\mathrm{Col}(V_{j,\ell^{\prime}})$, for $\ell^{\prime}\neq\ell$.
We then present the consistency results for $\widehat{r}_{j,U}$, whose proof is provided in Section D.2 of the supplementary material.
Theorem 3.2.
Under the assumptions of Theorem 3.1, further assume that
$\eta_{V}=\min_{j,\ell}\frac{1}{L-1}\sum_{\ell^{\prime}\neq\ell}(1-\left\lVert V_{j,\ell}^{\top}V_{j,\ell^{\prime}}\right\rVert_{2})\geq\eta_{0},$
for some constant $\eta_{0}>0$. For any $30<\delta_{0}<\eta_{0}/$ 3, we have
$\mathbb{P}\left(\widehat{r}_{j,U}=r_{j,U}\right)\geq 1-\widetilde{C}_{1}e^{-\widetilde{c}_{1}p},$
for some constants $\widetilde{c}_{1},\widetilde{C}_{1}>0$.
Theorem 3.2 demonstrates that, for a sufficiently small constant $\delta_{0}$, the estimator $\widehat{r}_{j,U}$ is consistent under the requirement $\eta_{V}\geq\eta_{0}$. This requirement ensures adequate heterogeneity among $V_{j,\ell}$ components. Indeed, the quantity $\eta_{V}=0$ if there exists a common principal vector across all $\mathrm{Col}(V_{j,\ell})$, as this implies $\left\lVert V_{j,\ell}^{\top}V_{j,\ell^{\prime}}\right\rVert_{2}=1$ for all $\ell^{\prime}\neq\ell$. Conversely, $\eta_{V}>0$ indicates that no principal vector is universally shared across all $\mathrm{Col}(V_{j,\ell})$ subspaces, thereby quantifying the heterogeneity among the individual components.
@@ 4 Knowledge Transfer in Distributed Tensor PCA
In this section, we explore the task of transferring knowledge from source locations to a target location within a heterogeneous setting. Knowledge Transfer seeks to enhance learning performance at a target site by leveraging insights from related source tasks. To illustrate our approach clearly, we concentrate on transferring knowledge between a single source dataset and a target dataset. The generalization to knowledge transfer across multiple tasks is provided in Section C of the supplementary material.
@@ 4.1 Transferred Tensor PCA
10:     Compute $\widehat{V}_{j,t}=\mathrm{svd}_{r_{j,V,t}}\Big{[}\big{(}I_{p_{j}}-\widehat{U}_{j}\widehat{U}_{j}^{\top}\big{)}\widetilde{M}_{j,t}\Big{]}$;
@@
$\mathcal{T}_{s}^{*}$ $=\mathcal{G}_{s}\times_{1}[U_{1}\;V_{1,s}]\times_{2}[U_{2}\;V_{2,s}]\cdots\times_{J}[U_{J}\;V_{J,s}],~{}U_{j}\in\mathbb{O}^{p_{j}\times r_{j,U}},V_{j,s}\in\mathbb{O}^{p_{j}\times r_{j,V,s}},\mathcal{G}_{s}\in\mathbb{R}^{r_{1,s}\times r_{2,s}\cdots\times r_{J,s}},$
$\mathcal{T}_{t}^{*}$ $=\mathcal{G}_{t}\times_{1}[U_{1}\;V_{1,t}]\times_{2}[U_{2}\;V_{2,t}]\cdots\times_{J}[U_{J}\;V_{J,t}],~{}U_{j}\in\mathbb{O}^{p_{j}\times r_{j,U}},V_{j,t}\in\mathbb{O}^{p_{j}\times r_{j,V,t}},\mathcal{G}_{t}\in\mathbb{R}^{r_{1,t}\times r_{2,t}\cdots\times r_{J,t}},$
with $U_{j}^{\top}V_{j,\ell}=0$ and $r_{j,\ell}=r_{j,U}+r_{j,V,\ell}$ for $j\in[J]$ and $\ell=s,t$. In other words, we assume the source and target tensors share a common top-$r_{j,U}$ singular subspace spanned by $U_{j}$, but either task can have different individual components $V_{j,\ell}$. The goal is to estimate the singular subspaces spanned by $U_{j}$ and $V_{j,t}$ of the target tensor $\mathcal{T}_{t}$. Meanwhile, we assume the noise $\mathcal{Z}_{\ell}$ has i.i.d. $\mathcal{N}(0,\sigma_{\ell}^{2})$ entries for $\ell=s,t$, where the noise levels $\sigma_{s}$ and $\sigma_{t}$ are allowed to be different. To achieve knowledge transfer between $\mathcal{T}_{s}$ and $\mathcal{T}_{t}$, we propose Algorithm 3, which is carefully designed for dealing with the heterogeneity in the transfer setting. Different from Algorithm 2 who treats all tensors equally, Algorithm 3 aggregates the local estimators $\widehat{U}_{j,s}$ and $\widehat{U}_{j,t}$ through a weighted average. The weights $w_{s}$ and $w_{t}$ are designed to optimally balance the contributions from the source and target tensors, respectively, accounting for the potential heterogeneity in their noise levels $\sigma_{s}^{2}$ and $\sigma_{t}^{2}$. The choice for the weights will be specified in the next section.
@@ 4.2 Theoretical Guarantee
Assume that there exist constants $C_{1},c_{1},C_{2}$ such that, with probability at least $1-C_{1}e^{-c_{1}p}$, $\big{\|}\widehat{U}_{j,\ell}^{(0)}\widehat{U}_{j,\ell}^{(0)\top}+\widehat{V}_{j,\ell}^{(0)}\widehat{V}_{j,\ell}^{(0)\top}-U_{j}U_{j}^{\top}-V_{j,\ell}V_{j,\ell}^{\top}\big{\|}_{2}\leq C_{2}\sqrt{p}\sigma_{\ell}\lambda_{\min}^{-1}$ for all $j\in[J],\ell\in\{s,t\}$. Assume $p_{j}\asymp p$ for all $j$, $r^{J-1}=O(p)$, $\kappa_{0}=O(1)$, $\kappa=O(1)$, and $\sqrt{pr}=o\big{(}\min(\Delta,\lambda_{\min})/\max(\sigma_{t},\sigma_{s})\big{)}$. There exists constants $\widetilde{C}_{1},\widetilde{c}_{1},\widetilde{C}_{2}$ such that
$\sup_{j}\left\lVert\widehat{U}_{j}\widehat{U}_{j}^{\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}$ $\leq\widetilde{C}_{2}\frac{\sqrt{pr}\sqrt{w_{s}^{2}\sigma_{s}^{2}+w_{t}^{2}\sigma_{t}^{2}}}{\Delta},$ (12)
$\sup_{j}\left\lVert\widehat{V}_{j,t}\widehat{V}_{j,t}^{\top}-V_{j,t}V_{j,t}^{\top}\right\rVert_{\rm F}$ $\leq\widetilde{C}_{2}\left(\frac{\sqrt{pr_{V}}}{\lambda_{\min}}\Big{(}\sigma_{t}+\sqrt{r}\sqrt{w_{s}^{2}\sigma^{2}_{s}+w_{t}^{2}\sigma_{t}^{2}}\Big{)}\right),$ (13)
with probability at least $1-\widetilde{C}_{1}e^{-\widetilde{c}_{1}p}$, where $\widehat{U}_{j},\widehat{V}_{j,t}$’s are the outputs of Algorithm 3.
@@
where $\overline{\sigma}^{2}=1/(\sigma_{s}^{-2}+\sigma_{t}^{-2})$. In practice, we can estimate $\sigma_{\ell}$ for $\ell=s,t$ by
$\widehat{\sigma}_{\ell}=\Big{\|}\mathcal{T}_{\ell}-\mathcal{T}_{\ell}\times_{1}\big{[}\widehat{U}_{1,\ell}\;\widehat{V}_{1,\ell}\big{]}\big{[}\widehat{U}_{1,\ell}\;\widehat{V}_{1,\ell}\big{]}^{\top}\times_{2}\cdots\times_{J}\big{[}\widehat{U}_{J,\ell}\;\widehat{V}_{J,\ell}\big{]}\big{[}\widehat{U}_{J,\ell}\;\widehat{V}_{J,\ell}\big{]}^{\top}\Big{\|}_{\rm F}/\sqrt{p_{1}p_{2}\cdots p_{J}},$ (14)
and estimate the optimal weights by $\widehat{w}_{s}=\frac{\widehat{\sigma}_{t}^{2}}{\widehat{\sigma}_{s}^{2}+\widehat{\sigma}_{t}^{2}}$ and $\widehat{w}_{t}=\frac{\widehat{\sigma}_{s}^{2}}{\widehat{\sigma}_{s}^{2}+\widehat{\sigma}_{t}^{2}}$.
In Theorem C.1 of the supplementary material, we generalize our knowledge transfer methodology and theoretical guarantee from the two-machine setting into an arbitrary number of machines $L$. Furthermore, we establish the statistical error rate for the estimators using data-adaptive weights $\widehat{w}_{\ell}$ in Theorem C.2.
@@ 5 Numerical Study
In this section, we conduct numerical studies to verify the theoretical properties and evaluate the empirical performance of our proposed distributed tensor PCA algorithms. Section 5.1 presents simulation studies for both the homogeneous and heterogeneous settings. Section 5.2 illustrates the performance of our algorithms on two real datasets of molecule structure.